# HandProxy: Expanding the Affordances of Speech Interfaces in Immersive Environments with a Virtual Proxy Hand

### v1

@@ 7. Discussion and Future Work
Here, we expand on our key study findings and discuss their implications for future work, including ambiguity handling, system transparency, additional support to bimanual interaction, and a unified accessibility API for hand interaction.
@@ 7.1. Handling Ambiguity in Interaction Commands
During the user study, the commands that were considered “ambiguous” by the system were often due to insufficient understanding of the environment and the target objects, including visual features, complex spatial referencing, and object affordance. Utilizing common sense and reasoning capabilities of LLMs, HandProxy aims to generalize across diverse ways a user may describe their interaction intention, with only a few tags and labels as metadata to capture the characteristics of supported gestures and target objects. In cases of ambiguity, the system sometimes took initiatives and filled in the necessary information. While this worked in some cases, it inevitably caused unexpected behavior (e.g., interpreting “turn the volume slider to maximum” as rotate volume knob to the right, rather than moving up). This highlights a design challenge that even with additional information about the environment, it is important to determine what can and should be automatically inferred by the system, and what should be prompted to users for clarification. An avenue for future work is to investigate how the system should balance automated inference and user clarification to consider both accuracy and user effort.
@@ 7.2. Supporting Multiple Levels of Interaction Control
Hand interaction commands could have varying levels of granularity. We summarized these variations into a spectrum of hand interaction commands in Figure 9, and showed the trade-offs in gesture reproducibility, ambiguity, user workload, and required system interpretation capability. At the high-granularity end of the spectrum, users have full precise control of the hand, including rotation and movement for the hand, finger, or specific joints. This allows users to reproduce almost any hand gesture and even fine-tune or customize gestures. While such commands tend to have low ambiguity, they impose a higher cognitive and physical workload on users due to the detailed input required. Conversely, at the low-granularity end of the spectrum, users can issue abstract, high-level interaction goals, leaving the system to determine how and what to do to achieve the goal. While this reduces user effort and simplifies command input, it comes with the trade-offs of reduced precise control, increased ambiguity, and a greater demand on the system’s interpretation capabilities.
@@
a spectrum from high to low of command granularity, gesture reproducibility, possible ambiguity, and user workload. From high to low granularity, command examples include ”move forward 10 units and bend all fingers 90 degrees, move forward stop and bend all fingers half way, move to the apple and grab it, grab the apple, and get me the red stuff.
@@ 7.3. Improving Transparency and Gesture Discoverability
During the user study, participants often struggled to identify the specific part of their commands caused issues when errors occurred. We believe the feedback system could be improved by providing more detailed information about how the system interprets commands. Specifically, it should include the interpreted action, the target, the preview of the decomposed hand movement, and should highlight the ambiguous parts and how system interpreted them. These could be achieved through a better-designed text feedback of the recognized command, such as to color-code the words in the command by the level of ambiguity and highlight the parts that could not be interpreted. Without overwhelm users with excessive feedback, such improvements could enhance the transparency of the system, help users effectively refine their inputs and improve their overall experience.
@@ 7.5. Towards an Accessible Interface for Hand Interactions
A key motivation of this work is to find an expressive, flexible alternative to hand interactions in cases of situational impairments, ability mismatches, or user’s varied preferences. While this work demonstrated how speech interface can be enhanced to achieve this goal, it brings up a broader question: can we define a unified control interface for possible hand interactions in the virtual environment — just like defining an API for 2D cursor controls, so that the interactions can be mapped to a broader range of input modality setup, potentially accommodating a wide range of user’s abilities and preferences?
To achieve this, it is important to create a comprehensive vocabulary for hand interactions in the virtual environment. Then, the direct mapping between the hand input and an arbitrary input modality could be simplified to the mapping to this shared vocabulary. This approach could enable mapping of hand input to other input setups, including multimodal configurations, that could be used to provide a more accessible way of interacting with the virtual environment.

### v2

@@ 7. Discussion and Future Work
Here, we expand on our key findings and discuss their implications for future work, including handling ambiguity, system transparency, supporting bimanual interaction, and creating a unified accessibility API for hand interaction.
@@ 7.1. Handling Ambiguity in Interaction Commands
During the user study, the commands that were considered “ambiguous” by the system were often due to insufficient understanding of the environment and the target objects, including visual features, complex spatial referencing, and object affordance. The current disambiguation design is mostly focusing on disambiguating duplicate objects. However, as discovered in the user study, additional ambiguity could come from the use of synonyms, or commands that have multiple interpretations. In cases of ambiguity, the system sometimes took initiatives and filled in the necessary information. While this worked in some cases, it inevitably caused unexpected behavior (e.g., interpreting “turn the volume slider to maximum” as rotate volume knob to the right, rather than moving up). This highlights a design challenge that even with additional information about the environment, it is important to determine what can and should be automatically inferred by the system, and what should be prompted to users for clarification. An avenue for future work is to investigate how the system should balance automated inference and user clarification to consider both accuracy and user effort.
@@ 7.2. Supporting Multiple Levels of Interaction Control
Hand interaction commands can vary in granularity along the spectrum shown in Figure 9, with trade-offs in gesture reproducibility, ambiguity, user workload, and system capability. At the high-granularity end of the spectrum, users have full, precise control of the hand. This allows users to reproduce almost any hand gesture and even fine-tune or customize gestures. While such commands are less ambiguous, they impose a higher cognitive and physical workload on users due to the detailed input required. Conversely, at the low-granularity end of the spectrum, users can issue abstract, high-level interaction goals, leaving the system to determine how and what to do to achieve the goal. While this reduces user effort and simplifies command input, it comes with the trade-offs of reduced precise control, increased ambiguity, and a greater demand on the system’s interpretation capabilities.
@@
A spectrum from high to low of command granularity, gesture reproducibility, possible ambiguity, and user workload. From high to low granularity, command examples include ”move forward 10 units and bend all fingers 90 degrees, move forward stop and bend all fingers half way, move to the apple and grab it, grab the apple, and get me the red stuff. An ideal range of supported granularity should not be too specific and also not too generic, roughly 10% to 95% along the spectrum.
@@ 7.3. Improving Transparency and Gesture Discoverability
During the user study, participants often struggled to identify the specific part of their commands caused issues when errors occurred. We believe the feedback system could be improved by providing more details about how the system interprets commands, and should highlight the ambiguous parts and how the system interpreted them. These could be achieved through a better-designed text feedback of the recognized command, such as to color-code the words in the command by the level of ambiguity and highlight the parts that could not be interpreted. Without overwhelm users with excessive feedback, such improvements could enhance the transparency of the system, help users effectively refine their inputs and improve their overall experience.
@@ 7.4. Supporting Bimanual Interactions
7.5. Limitation on Applicable Use Cases
While the proposed approach is designed to generalize across a wide range of interaction scenarios, it may not always offer substantial advantages over alternative methods. For more direct and less hand-dependent interactions – such as basic system operations (e.g., power on/off) or standard UI controls (e.g., closing windows) – a traditional speech interface with direct command mapping may be more effective. Interactions requiring high precision or complexity, such as fine-grained rotations in training or simulation-based XR applications, may pose challenges when relying solely on the proposed speech interface. Furthermore, although our approach leverages commonly used gestures across various immersive applications, the current implementation does not readily accommodate applications that heavily depend on customized, non-standard hand gestures. For example, the spell-casting game Drakheir Hands of Wizard [^4] https://www.meta.com/experiences/drakheir-hands-of-wizard requires specialized gesture inputs that are not easily replicated within the existing system. These observations suggest that further enhancements are needed to support a broader range of use cases, particularly those involving diverse gesture types, varying levels of precision, and additional control requirements.
As an initial exploration of the proxied interaction paradigm for speech interfaces, HandProxy shows the potential of using virtual hands to broaden interaction possibilities. Future work could address current limitations through the integration of multimodal input strategies that adaptively optimize interaction based on task type and complexity. Additional improvements might include support for runtime gesture recording to enable customizable gesture macros, as well as integration with system-level APIs to facilitate cross-application control.
@@ 7.6. Towards an Accessible Interface for Hand Interactions
A key motivation of this work is to find an expressive, flexible alternative to hand interactions in cases of situational impairments, ability mismatches, or user’s varied preferences. While this work demonstrated how speech interfaces can be enhanced to achieve this goal, it brings up a broader question: can we define a unified control interface for possible hand interactions in the virtual environment — just like defining an API for 2D cursor controls, so that the interactions can be mapped to a broader range of input modality setups, potentially accommodating a wide range of user’s abilities and preferences?
To achieve this, it is important to create a comprehensive, well-defined set of vocabularies for hand interactions in the virtual environment. Then, the direct mapping between the hand input and an arbitrary input modality could be simplified to the mapping to this shared vocabulary. This approach could enable flexible and extensible mapping of hand input to other input setups, including multimodal configurations, that could be used to provide a more accessible way of interacting with the virtual environment.
@@ 8. Conclusion
Appendix A Example Gesture Metadata File
"name": "cut",
"data_format": "unified",
"data_source": "leap_motion",
"num_hands": 1,
"right_hand_data_file": "cut.txt",
"left_hand_data_file": null,
"is_hold_at_peak": false,
"is_static": false,
"interacting_frame": 81,
"interacting_joint": ["pinky_mcp"],
"segments": [
"name": "preparation",
"start_frame": 0,
"end_frame": 58
"name": "stroke",
"start_frame": 58,
"end_frame": 123
"name": "retraction",
"start_frame": 123,
"end_frame": 200